Find the amount that results from the given investment. $500 invested at 9% compounded quarterly after a period of 2 years After 2 years, the investment results in $ (Round to the nearest cent as needed.)
Answer #1

A=P\left(1+\frac{r}{n}\right)^{nt}

p=500

r=9%=0.09

t=2 year

n=4 (compound quartly)

A=500\left(1+\frac{0.09}{4}\right)^{4\cdot 2}

A=500\left(1.0225\right)^8

A=500\cdot \:1.19483

{\color{Red} A=597.41557{\color{Green}}}